pooja195 (pooja195):

A rectangle has perimeter 18. If each side of the rectangle is doubled, the perimeter of the new rectangle is 36. What were the dimensions of the original rectangle?

OpenStudy (anonymous):

Whatcha thank boi?

pooja195 (pooja195):

GIRL xD i dunno ;-;

OpenStudy (anonymous):

think*

OpenStudy (anonymous):

ok, what does it mean to double?

pooja195 (pooja195):

2

OpenStudy (anonymous):

ok, 18*2=??

pooja195 (pooja195):

36?

OpenStudy (anonymous):

mmm-k, iws dis a test?

pooja195 (pooja195):

LOL xD nawh i dont have the guts to post one :P btw ru gonna help? im lost

pooja195 (pooja195):

that bit of info is already there?

OpenStudy (anonymous):

ok, Let l be the length and w the width of the original rectangle. Then\[2l+2w=18\]

pooja195 (pooja195):

huh????

OpenStudy (anonymous):

\[2(2l)+2(2x)=36\]

pooja195 (pooja195):

where are you getting this from? im confused

OpenStudy (anonymous):

thats not the answer?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

i is too

pooja195 (pooja195):

@Zarkon ?

OpenStudy (zarkon):

did you make this problem up?

pooja195 (pooja195):

nope

OpenStudy (zarkon):

well...it does not have a unique solution

pooja195 (pooja195):

hmm...thats weird i'll ask my teacher about this one thanks :)

OpenStudy (zarkon):

I can give you two examples rectangle 1...L=5,W=4, 5+5+4+4=18, 2*5+2*5+2*4+2*4=36 rectangle 2 ...L=7,W=2, 7+7+2+2=18, 2*7+2*7+2*2+2*2=36 both work

pooja195 (pooja195):

O_O wuw so the idea of the problem is to do a bunch of math and try to get 36 using numbers?

OpenStudy (zarkon):

no...not a lot of math...I can give you all the possibilities really easily L=x W=9-x then\( L+L+W+W=x+x+9-x+9-x=18\) \(2L+2L+2W+2W=2x+2x+(18-2x)+(18-2x)=36\) pick for favorite value for x (0<x<9)